# BillBoard Splatting (BBSplat): Learnable Textured Primitives for Novel View Synthesis

### v1

@@ Abstract
We present billboard Splatting (BBSplat) - a novel approach for 3D scene representation based on textured geometric primitives. BBSplat represents the scene as a set of optimizable textured planar primitives with learnable RGB textures and alpha-maps to control their shape. BBSplat primitives can be used in any Gaussian Splatting pipeline as drop-in replacements for Gaussians. Our method’s qualitative and quantitative improvements over 3D and 2D Gaussians are most noticeable when fewer primitives are used, when BBSplat achieves over 1200 FPS. Our novel regularization term encourages textures to have a sparser structure, unlocking an efficient compression that leads to a reduction in storage space of the model. Our experiments show the efficiency of BBSplat on standard datasets of real indoor and outdoor scenes such as Tanks&Temples, DTU, and Mip-NeRF-360. We demonstrate improvements on PSNR, SSIM, and LPIPS metrics compared to the state-of-the-art, especially for the case when fewer primitives are used, which, on the other hand, leads to up to inference speed improvement for the same rendering quality.
@@ 1 Introduction
Indeed, NeRF-based methods [34, 5, 4, 35] still achieve the best NVS quality for challenging real-world captures by using implicit scene representations such as the weights of an MLP. However, image rendering with a NeRF is less efficient as it requires repeated MLP inferences to predict colors along camera rays. An alternative to neural rendering, 3D Gaussian Splatting (3DGS) [23] used faster rendering based on projecting explicit primitives on the screen surface while preserving high-quality NVS. In practice, 3DGS uses Gaussian-distributed radiance around explicit 3D scene points as primitives.
@@
Our proposed geometric primitives for NVS take inspiration from the classic billboards used for extreme 3D model simplification [9] by replacing mesh during a greedy optimization process. A 3D scene can be efficiently rendered by using a “billboard cloud” of given textured planar primitives with alpha channels. Using billboards, we can efficiently model planar surfaces, such as a painting on a wall or scene background, while dramatically reducing the number of geometric primitives required, up to an order of magnitude with respect to 3DGS/2DGS (check LABEL:fig:shapes for a comparison). In combination with efficient texture sampling implemented on the GPU, this leads to inference speed improvements without a drop in the rendering quality (Section 4.2).
@@
To ensure rendering efficiency, we implemented texture sampling and a back-propagation process in CUDA. In this way, the use of textures does not result in overtime compared to the 2DGS Gaussians’ rasterization process. To tackle the challenge of storing all billboard textures, we compress them by representing each texture as sparse offsets from colors predicted by spherical harmonics [2] and by further quantizing them to 8 bits. Then, we can efficiently utilize dictionary-based compression algorithms [41, 11] for quantized textures.
@@ 2 Related work
We noted that existing approaches use a tremendous number of Gaussians to represent the scene, which affects rendering speed. Billboard clouds [9] propose an extreme 3D model simplification by replacing it with a set of textured planes. This representation is widely used in computer graphics for real-time forest rendering [3, 14, 27] and found implementation in such tasks as automatic paper slice-form design [28, 33, 39] and LOD control [17]. In this work, we take inspiration from the billboard clouds technique [9] and propose arbitrarily shaped textured primitives as an extension of the 2DGS approach, resulting in faster and more qualitative rendering of real scenes reconstructed from the set of photographs.
@@ 3.1 Preliminaries: 2D Gaussian Splatting
Recently proposed 2D Gaussian splitting (2DGS) [19] uses 2D flat ellipses (in contrast with 3D ellipsoids for 3DGS) orientated along the object’s surface to represent the scene. These ellipses have Gaussian distributed transparency and are parametrized as which corresponds to position, 2D scale, quaternion rotation, global opacity, and spherical harmonics of the primitive as in Fig. 1(b1). Here spherical harmonics [2, 12] define view-dependent color. In practice, to calculate ellipses orientation, 2DGS uses local tangent planes defined with transformation matrix as:
@@ 3.3 Training
To train billboard splatting representation of the scene, we initialize them with sparse point cloud obtained using SfM [40]. Optionally, we add 10K evenly distributed points on the circumscribing sphere using the Fibonacci algorithm [16] to represent the sky and far away objects. We use only photometric losses to fit a scene as in Gaussian splatting pipelines. For textures, we propose regularizations to avoid their overfitting and get a more sparse structure that reduces storage costs. In this section, we describe the losses and regularizations we used.
Photometric losses. We utilize photometric losses to train our billboard representation. Namely we apply and structure similarity losses:
@@
Regularizations. Due to the large number of parameters, textures tend to overfit training images, which results in noisy rendering output for novel views. To alleviate this issue, we propose a simple yet efficient regularization in which we push billboards with low impact on the renders (i.e. small or far billboards affecting few pixels) to have Gaussian distributed transparency.
@@
Alpha-blending values are calculated similarly to 2DGS and take into account overlapping planes at pixel . We use the maximum impact threshold to regularize only clearly visible billboards. The choice of the threshold is discussed in the supplementary materials. By ignoring billboards with , we do not regularize invisible in the frame primitives, avoiding over-regularization in underrepresented areas. The proposed criterion is more efficient than the one utilizing projected radii as in 2DGS since they only take hitting frustum into account and ignore overlapping.
@@ 4.1 Implementation
For training, we use the Adam [25] optimizer with hyper-parameters recommended for the Gaussian splitting pipelines [23]. We set RGB and textures size . resulting in textures to get the best trade-off between quality and memory requirements. To optimize RGB-texture we set learning rate and for -texture we set . Additionally, we found it beneficial for convergence to increase spherical harmonics learning rate to .
@@ 4.2 Results and Evaluation
We compared with previous state-of-the-art methods on real-world scenes using standart datasets. We evaluated our method on 4 indoor scenes from the Mip-NeRF-360 dataset [5] and 5 outdoor scenes from the Tanks&Temples dataset [26]. Additionally, we evaluated our method on 15 scenes from DTU dataset [22]. Overall, we provide experiments for 24 real scenes from three datasets.
@@
Qualitative comparison. We provide qualitative results in Figure 4 for indoor and outdoor scenes from different datasets. We show novel views for basic 3DGS and 2DGS representations and 3DGS-MCMC as they are the best scoring competitors. While the quality of the central object of the scene remains similar for all methods, in most scenarios BBSplat represents background objects with more details. For Tanks&Temples and Mip-NeRF-360 dataset, we reached quality on par with state-of-the-art while providing faster inference speed and higher level of detail. For the DTU dataset, we also demonstrate better handling of challenging shadow effects. For more results, please refer to supplementary materials.
4.3 Ablation study.
@@
In Table 4, we provide an ablation study of different aspects of our approach. Namely the use of RGB and alpha textures, texture regularization term, and compression post-processing.
The use of both alpha and RGB textures is essential for rendering quality. Excluding significantly increases LPIPS, while excluding reduces PSNR and SSIM. Texture regularization term preserves over-fitting and leads to additional metrics improvement. It also encourages sparsity in textures and reduces storage costs in combination with compression. The use of compression leads to a slight metric reduction but allows to significantly reduce storage space requirements.
Therefore, all parts of the proposed approach are essential to get the best result. However, we note that metrics could be additionally improved by disabling compression.

### v2

@@ Abstract
We present billboard Splatting (BBSplat) - a novel approach for 3D scene representation based on textured geometric primitives. BBSplat represents the scene as a set of optimizable textured planar primitives with learnable RGB textures and alpha-maps to control their shape. BBSplat primitives can be used in any Gaussian Splatting pipeline as drop-in replacements for Gaussians. Our method’s qualitative and quantitative improvements over 3D and 2D Gaussians are most noticeable when fewer primitives are used, when BBSplat achieves over 1200 FPS. Our novel regularization term encourages textures to have a sparser structure, unlocking an efficient compression that leads to a reduction in storage space of the model. Our experiments show the efficiency of BBSplat on standard datasets of real indoor and outdoor scenes such as Tanks&Temples, DTU, and Mip-NeRF-360. We demonstrate improvements on PSNR, SSIM, and LPIPS metrics compared to the state-of-the-art, especially for the case when fewer primitives are used, which, on the other hand, leads to up to inference speed improvement for the same rendering quality. Project page: david-svitov.github.io/BBSplat_project_page.
@@ 1 Introduction
Indeed, NeRF-based methods [34, 5, 4, 35] still achieve the best NVS quality for challenging real-world captures by using implicit scene representations such as the weights of an MLP. However, image rendering with a NeRF is less efficient as it requires repeated MLP inferences to predict colors along camera rays. An alternative to neural rendering, 3D Gaussian Splatting (3DGS) [23] uses faster rendering based on projecting explicit primitives on the screen surface while preserving high-quality NVS. In practice, 3DGS uses Gaussian-distributed radiance around explicit 3D scene points as primitives.
@@
Our proposed geometric primitives for NVS take inspiration from the classic billboards used for extreme 3D model simplification [9] by replacing mesh during a greedy optimization process. A 3D scene can be efficiently rendered by using a “billboard cloud” of given textured planar primitives with alpha channels. Using billboards, we can efficiently model planar surfaces, such as a painting on a wall or scene background, while dramatically reducing the number of geometric primitives required, up to an order of magnitude with respect to 3DGS/2DGS (check LABEL:fig:shapes for a comparison). In combination with efficient texture sampling implemented on the GPU, this leads to inference speed improvement without a drop in the rendering quality.
@@
To ensure rendering efficiency, we implemented texture sampling and a back-propagation process in CUDA. In this way, the use of textures does not result in overtime compared to the 2DGS Gaussians’ rasterization process. To tackle the challenge of storing all billboard textures, we compress them by representing each texture as sparse offsets from colors calculated by spherical harmonics [2] and by further quantizing them to 8 bits. Then, we can efficiently utilize dictionary-based compression algorithms [41, 11] for quantized textures.
@@ 2 Related work
We noted that existing approaches use a tremendous number of Gaussians to represent the scene, which affects rendering speed. Billboard clouds [9] propose an extreme 3D model simplification by replacing it with a set of textured planes. This representation is widely used in computer graphics for real-time forest rendering [3, 14, 27] and found implementation in such tasks as automatic paper slice-form design [28, 33, 39] and level of detail control [17]. In this work, we take inspiration from the billboard clouds technique [9] and propose arbitrarily shaped textured primitives as an extension of the 2DGS approach, resulting in faster and more qualitative rendering of real scenes reconstructed from the set of photographs.
@@ 3.1 Preliminaries: 2D Gaussian Splatting
Recently proposed 2D Gaussian splatting (2DGS) [19] uses 2D flat ellipses (in contrast with 3D ellipsoids for 3DGS) orientated along the object’s surface to represent the scene. These ellipses have Gaussian distributed transparency and are parametrized as which corresponds to position, 2D scale, quaternion rotation, global opacity, and spherical harmonics of the primitive as in Fig. 1(b1). Here spherical harmonics [2, 12] define view-dependent color. In practice, to calculate ellipses orientation, 2DGS uses local tangent planes defined with transformation matrix as:
@@ 3.3 Training
To train billboard splatting representation of the scene, we initialize them with sparse point cloud obtained using SfM [40]. Optionally, we add 10K evenly distributed points on the circumscribing sphere using the Fibonacci algorithm [16] to represent the sky and far away objects. We use only photometric losses to fit a scene as in Gaussian splatting pipelines. For textures, we propose regularizations to avoid their overfitting and get a more sparse structure that reduces storage costs. In this section, we describe the losses and regularizations we use.
Photometric losses. We utilize photometric losses to train our billboard representation. Namely we apply and structure similarity D-SSIM losses:
@@
Regularizations. Due to the large number of parameters, textures tend to overfit training images, which results in noisy rendering output for novel views. To alleviate this issue, we propose a simple yet efficient regularization in which we push billboards with low impact on the renderings (i.e. small or far billboards affecting few pixels) to have Gaussian distributed transparency.
@@
Alpha-blending values are calculated similarly to 2DGS and take into account overlapping planes at pixel . We use the maximum impact threshold to regularize only clearly visible billboards. The choice of the threshold is discussed in the Sec. 6.2 of supp. mat.. By ignoring billboards with , we do not regularize invisible in the frame primitives, avoiding over-regularization in underrepresented areas. The proposed criterion is more efficient than the one utilizing projected radii as in 2DGS since they only take hitting frustum into account and ignore overlapping.
@@ 4.1 Implementation
For training, we use the Adam [25] optimizer with hyper-parameters recommended for the Gaussian splatting pipelines [23]. We set RGB and textures size . resulting in textures to get the best trade-off between quality and memory requirements. To optimize RGB-texture we set learning rate and for -texture we set . Additionally, we found it beneficial for convergence to increase spherical harmonics learning rate to .
@@ 4.2 Results and Evaluation
We compared with previous state-of-the-art methods on real-world scenes using standard datasets. We evaluated our method on 4 indoor scenes from the Mip-NeRF-360 dataset [5] and 5 outdoor scenes from the Tanks&Temples dataset [26]. Additionally, we evaluated our method on 15 scenes from DTU dataset [22]. Overall, we provide experiments for 24 real scenes from three datasets.
@@
Qualitative comparison. We provide qualitative results in Figure 4 for indoor and outdoor scenes from different datasets. We show novel views for basic 3DGS and 2DGS representations and 3DGS-MCMC as they are the best scoring competitors. While the quality of the central object of the scene remains similar for all methods, in most scenarios BBSplat represents background objects with more details. For Tanks&Temples and Mip-NeRF-360 dataset, we reached quality on par with state-of-the-art while providing faster inference speed and higher level of detail. For the DTU dataset, we also demonstrate better handling of challenging shadow effects. For more results, please refer to Sec. 7.1 of supp. mat..
4.3 Ablation study
@@
In Table 4, we provide an ablation study of different aspects of our approach. Namely the use of RGB and alpha textures, texture regularization, and compression post-processing.
The use of both alpha and RGB textures is essential for rendering quality. Excluding significantly increases LPIPS, while excluding reduces PSNR and SSIM. Texture regularization term preserves over-fitting and leads to additional metrics improvement. It also encourages sparsity in textures and reduces storage costs in combination with compression. The use of compression leads to a slight metric reduction but allows to significantly reduce storage space requirements. Therefore, all parts of the proposed approach are essential to get the best result. However, we note that metrics could be additionally improved by disabling compression.
@@ Acknowledgements.
Supplementary Material
| 10 | 100 | 500 | 1000 | 1500 | |
| PSNR | 25.95 | 25.96 | 25.99 | 25.92 | 25.92 |
| SSIM | 0.8800 | 0.8798 | 0.8798 | 0.8798 | 0.8795 |
| LPIPS | 0.1365 | 0.1375 | 0.1379 | 0.1392 | 0.1396 |
| Size(MB) | 139 | 122 | 102 | 92 | 87 |
| FPS | 995 | 1033 | 1067 | 1081 | 985 |
| PSNR | 26.04 | 25.99 | 25.28 | 23.34 |
| SSIM | 0.8807 | 0.8798 | 0.8662 | 0.8000 |
| LPIPS | 0.1393 | 0.1379 | 0.1505 | 0.2382 |
| Size(MB) | 207 | 102 | 82 | 68 |
| FPS | 618 | 1067 | 1211 | 1267 |
6 Implementation details
In this section, we provide more details on the method’s implementation. In the following subsections, we describe the implementation of backpropagation for the bilinear texture sampling with billboard position adjustment and discuss the choice of hyper-parameters to train billboard splatting.
6.1 Texture sampling
We leverage bilinear sampling to get texture values from and in the point of a billboard. Our implementation utilizes GPU programmed with CUDA to accelerate sampling and loss backpropagation to the texture. We based our implementation on the PyTorch[38] CUDA code for bilinear sampling and integrated it into 2DGS CUDA implementation as described below.
In Algorithm 1, we provide pseudo-code for the bilinear sampling gradients backpropagation. The algorithm accepts the gradient value that should be distributed among neighboring texels, coordinate of the billboard point where the gradient is calculated, corresponding source texture, and output tensor to store texture gradients. The function redistributes and stores weighted values of the gradient in four neighboring texels and calculate the gradient with respect to point position.
In Algorithm 2, we demonstrate how we prepare input parameters for Algorithm 1 to get values based on RGB texture and alpha map. Calculated value is then used to adjust billboard position and rotation. The proposed implementation can be easily integrated with the 2DGS approach, which we encapsulate here as subprogram . The function encapsulates gradients calculation for transformation matrix encoding translation and rotation of the plane.
6.2 Hyper-parameters
We apply learning rate exponential decay for billboards position from to with multiplier. To optimize spherical harmonics and scales we set while using smaller learning rate for rotation .
We select threshold value in Equation 9 based on the metrics provided in Table 5. The threshold is used to determine the minimum billboard visibility from which we start regularizing its textures. We set as it provides the best PSNR value and achieves the best trade-off between other metrics and storage space. Over-regularization results in more billboards having Gaussian distributed transparency, which leads to more compact storing of them but reduces NVS quality. Under-regularization allows to achieve higher SSIM and LPIPS at the cost of storage space and inference speed.
Table 6 shows metrics for different texture sizes and number of billboards. We select texture size to be the power of two (8, 16, 32, 64) as it is more suitable for processing with GPU, and set the corresponding number of billboards to result in 40.96M parameters for one texture channel to not exceed GPU capacity. While and both provide suitable metrics, we chose as slightly more accurate. As an alternative, can be used to achieve faster inference.
6.3 Initialization
In scenarios when we need fewer SfM points to initialize a small number of billboards, we subsample them with an iterative farthest point sampling strategy. For outdoor scenes, we optionally add 10K evenly distributed points on the large sphere using the Fibonacci algorithm [16] to represent the sky and far away objects. We select sphere radius as the distance from the scene center to the furthest SfM point to guarantee that all SfM points are included in the sphere. For scenarios when we use 20K points and less, we reduce the number of sphere points to 2K, and completely disable it when we use less than 10K points. In Figure 5, we provide a visualization of such a sphere.
7 Additional details on experiments and results
In this section, we provide more qualitative results for our method along with per-scene quantitative metrics for averaged values reported in the paper.
7.1 Visual results
In Figure 7, we provide an additional visual comparison of BBSplat with state-of-the-art NVS methods. Our renderings demonstrate exceeding quality for similar or higher FPS.
Figure 6 demonstrates BBSplat training progress for 500, 5000, 15000, and 30000 iterations, showing the change of transparency from Gaussian distribution to arbitrary shapes.
7.2 Detailed results
In Table 7, we report metrics for all scenes. For competitors, we provide metrics for the maximum recommended number of Gaussians, and for ours with 300K billboards. The averaged values for these three datasets are reported in the main paper. For examples of renderings from different angles, see the video in the supp. mat..
| 3DGS [23] | 3DGS-MCMC [24] | 2DGS† [19] | BBSplat (Ours) | ||
| PSNR / SSIM / LPIPS | PSNR / SSIM / LPIPS | PSNR / SSIM / LPIPS | PSNR / SSIM / LPIPS | ||
| Tanks&Temples | Train | 22.57 / 0.8248 / 0.1968 | 23.26 / 0.8531 / 0.1735 | 22.03 / 0.8146 / 0.2144 | 23.04 / 0.8336 / 0.1696 |
| Truck | 26.04 / 0.8851 / 0.1489 | 27.23 / 0.9024 / 0.1230 | 26.07 / 0.8814 / 0.1612 | 26.06 / 0.8800 / 0.1336 | |
| Francis | 28.00 / 0.9127 / 0.2349 | 29.24 / 0.9217 / 0.2275 | 28.23 / 0.9118 / 0.2424 | 30.08 / 0.9231 / 0.1960 | |
| Horse | 25.26 / 0.9054 / 0.1272 | 27.53 / 0.9227 / 0.1108 | 25.70 / 0.9044 / 0.1321 | 25.85 / 0.9027 / 0.1311 | |
| Lighthouse | 21.99 / 0.8482 / 0.2061 | 23.71 / 0.8701 / 0.1846 | 22.70 / 0.8556 / 0.2061 | 23.85 / 0.8627 / 0.1822 | |
| Mip-NeRF-360 | Room | 32.31 / 0.9401 / 0.1295 | 33.19 / 0.9507 / 0.1075 | 32.33 / 0.9370 / 0.1365 | 32.30 / 0.9454 / 0.1062 |
| Kitchen | 32.12 / 0.9370 / 0.1114 | 32.93 / 0.9429 / 0.1056 | 31.50 / 0.9303 / 0.1248 | 32.00 / 0.9365 / 0.1084 | |
| Bonsai | 32.26 / 0.9495 / 0.0977 | 32.77 / 0.9554 / 0.0861 | 32.03 / 0.9460 / 0.1055 | 31.85 / 0.9464 / 0.0940 | |
| Counter | 29.84 / 0.9210 / 0.1191 | 30.30 / 0.9303 / 0.1059 | 29.60 / 0.9142 / 0.1302 | 29.88 / 0.9201 / 0.1091 | |
| DTU | Scan24 | 25.64 / 0.9399 / 0.0780 | 29.28 / 0.9531 / 0.0605 | 26.59 / 0.9402 / 0.0889 | 26.89 / 0.9332 / 0.0914 |
| Scan37 | 25.81 / 0.9285 / 0.0973 | 27.58 / 0.9448 / 0.0757 | 25.81 / 0.9341 / 0.0870 | 24.73 / 0.9165 / 0.1019 | |
| Scan40 | 24.73 / 0.9278 / 0.1126 | 29.35 / 0.9528 / 0.0776 | 26.04 / 0.9218 / 0.1165 | 26.92 / 0.9323 / 0.0992 | |
| Scan55 | 28.48 / 0.9199 / 0.1750 | 30.66 / 0.9470 / 0.1093 | 28.59 / 0.9187 / 0.1834 | 29.50 / 0.9322 / 0.1331 | |
| Scan63 | 30.48 / 0.9677 / 0.0577 | 32.99 / 0.9767 / 0.0434 | 31.62 / 0.9671 / 0.0567 | 29.51 / 0.9528 / 0.0661 | |
| Scan65 | 27.69 / 0.9131 / 0.1913 | 30.65 / 0.9417 / 0.1332 | 29.30 / 0.9191 / 0.1982 | 29.59 / 0.9278 / 0.1476 | |
| Scan69 | 26.88 / 0.9142 / 0.1564 | 27.85 / 0.9306 / 0.1266 | 27.26 / 0.9154 / 0.1624 | 26.79 / 0.9168 / 0.1390 | |
| Scan83 | 29.64 / 0.9397 / 0.1876 | 30.59 / 0.9468 / 0.1670 | 29.69 / 0.9336 / 0.1975 | 30.08 / 0.9409 / 0.1831 | |
| Scan97 | 26.84 / 0.9180 / 0.1670 | 27.68 / 0.9246 / 0.1588 | 27.03 / 0.9172 / 0.1742 | 28.52 / 0.9272 / 0.1600 | |
| Scan105 | 30.23 / 0.9414 / 0.1511 | 32.26 / 0.9507 / 0.1381 | 30.06 / 0.9364 / 0.1659 | 31.23 / 0.9465 / 0.1479 | |
| Scan106 | 32.93 / 0.9443 / 0.1828 | 35.33 / 0.9628 / 0.1133 | 33.89 / 0.9469 / 0.1851 | 34.14 / 0.9502 / 0.1764 | |
| Scan110 | 31.67 / 0.9396 / 0.1945 | 33.49 / 0.9553 / 0.1530 | 31.88 / 0.9391 / 0.2011 | 32.52 / 0.9433 / 0.1812 | |
| Scan114 | 29.76 / 0.9266 / 0.1861 | 31.50 / 0.9518 / 0.1091 | 31.26 / 0.9366 / 0.1792 | 31.28 / 0.9383 / 0.1703 | |
| Scan118 | 33.40 / 0.9376 / 0.2074 | 36.00 / 0.9647 / 0.1142 | 34.40 / 0.9442 / 0.2063 | 35.37 / 0.9503 / 0.1813 | |
| Scan122 | 33.58 / 0.9436 / 0.2151 | 35.88 / 0.9626 / 0.1343 | 34.42 / 0.9435 / 0.2161 | 34.34 / 0.9487 / 0.1943 |